Adrielle17:

How many miles are presented in 10.5 grams of NaCl

xmoseje:

Mr of NaCl = 58.5 Mole = mass÷ Mr.. Mole od NaCl = 10.5 ÷ 58.5 = 0.179

xmoseje:

@Adrielle17